# Quantum mechanical dataset of 836k neutral closed shell molecules with upto 5 heavy atoms from CNOFSiPSClBr

### v4

@@ Methods
To further assess the chemical diversity we trained and tested ML models for the task of predicting atomization energies within VQM24 and compare it to ML results obtained for the commonly used QM9 benchmark. Fig. 4 shows learning curves (model prediction error as a function of training set size) of atomization energies on both datasets. All ML models use kernel-ridge regression (KRR) 79; 80 along with a few commonly used atomic representations. Since QM9 does not contain conformational isomers, Fig. 4A shows learning curves of atomization energies considering only the subset of the 258k lowest energy conformers (unique constitutional isomers) from VQM24. While the range of atomization energies covered by VQM24 (1545 kcal/mol) is smaller than QM9 (2427 kcal/mol), evidently they are more challenging to learn as all ML models show upto $\sim$8 times larger mean errors than on QM9 for the same training set size. This is likely due to the much larger chemical diversity of VQM24 which should make it a more challenging benchmark for the training and testing of ML models of the various chemically relevant physical properties reported.
@@
Using the best KRR model from Fig. 4 (cMBDF 57; 56), we performed a prediction error analysis to detect outliers within the dataset. We trained KRR/cMBDF based ML models on 200k molecules (4 disjointed training sets) to make predictions on the remaining dataset. This was done on all the 784,875 equilibrium geometries (including conformers) reported in the dataset. The resulting distribution of errors on the entire dataset is shown in fig. 5. For each molecule we used the prediction with the smallest error out of the 3 ML models that were not trained on it. The mean absolute error for the 784,875 predictions was obtained to be 0.75 kcal/mol with a standard deviation of 1.55 kcal/mol. The largest obtained error was 167.3 kcal/mol and the 25 largest outliers show a mean absolute error of 85.9 kcal/mol. These molecules are shown as insets in fig. 5.
Usage Notes
The reported .npz files available at https://doi.org/10.5281/zenodo.11164951 can be readily accessed using the Numpy 81 library in Python. DFT properties for all 784,875 conformers in local minima; 258,242 constitutional isomers (most stable conformer) and 51,072 saddle point structures are available in the DFT_all.npz, DFT_uniques.npz and DFT_saddles.npz files respectively. DMC data for 10,793 constitutional isomers is available in the DMC.npz file. All molecules are ordered in the same way across every array. Keys for accessing each property are tabulated in the Key column of tables  2 and  3. Usage example :
import numpy as np
data = np.load(’DFT_all.npz’, allow_pickle=True)
print(data.files) #see a list of all properties
key = ’freqs’
property = data[key] #DFT vibrational frequencies
of all molecules

### v5

@@ III Validation and analysis
Figure 6: Atomization energy learning curves on the subset of 258k unique constitutional isomers from VQM24 and the QM9 dataset. Solid lines indicate ML models employing Kernel Ridge Regression (KRR) while dashed lines indicate Graph Neural Networks (GNN). Representations used alongside KRR models are : Coulomb matrix 76 (CM), atom-centered symmetry functions 77 (ACSF), many-body tensor representation 78 (LMBTR), Faber-Christensen-Huang-Lilienfeld 19 79; 80 (FCHL19), convolutional many-body distribution functionals 81; 82 (cMBDF/MBDF). GNNs employ the equivariant PaiNN 83 and invariant SchNet 84 architectures. Test set size in both cases was 10,000 randomly-selected molecules. Plots show average of 5 such runs.
To further assess the chemical diversity we trained and tested ML models for the task of predicting atomization energies within VQM24 and compare it to ML results obtained for the commonly used QM9 benchmark. Fig. 6 presents the learning curves—i.e. the prediction error on atomization energies as a function of training‐set size—for both datasets. We compare several kernel ridge regression (KRR) 85; 86 models using different atomic representations, alongside invariant (SchNet 84) and equivariant (PaiNN 83) message‐passing graph neural networks (GNNs). The KRR and GNN models were trained and deployed using the QMLcode 87 and Schnetpack 88 libraries respectively. Atomic gaussian kernel was used alongside all KRR models and the hyper-parameters (length-scale $l$, regularizer $\lambda$) were optimized via grid-search. Logarithmic grids of $\left[0.1(2^{n})~{}\forall~{}n~{}\in\{0,14\}\right]$ and $\left[10^{-3n}~{}\forall~{}n~{}\in\{1,4\}\right]$ were employed for $l$ and $\lambda$ respectively. Optimizations were performed via 5-fold cross-validation. For both GNN models we employed the same hyper-parameters as in the PaiNN 83 paper. We note here that we used 128 atomic basis functions with both PaiNN and SchNet leading to a total parameter count of 589k for both models. All models were trained for 1000 epochs using the Adam optimizer 89 with a learning rate of $10^{-4}$. Scripts for training and using these KRR and GNN models are available in the GitHub repository specified in the Code Availability section below.
@@
Since QM9 does not contain conformational isomers, Fig. 6B shows learning curves of atomization energies considering only the subset of the 258k lowest energy conformers (unique constitutional isomers) from VQM24 to be directly comparable. While the range of atomization energies covered by VQM24 (1545 kcal/mol) is smaller than QM9 (2427 kcal/mol), evidently they are more challenging to learn as all ML models show upto $\sim$8 times larger mean errors than on QM9 for the same training set size. This is likely due to the much larger chemical diversity of VQM24 which should make it a more challenging benchmark for the training and testing of ML models of the various chemically relevant physical properties reported.
@@
Using the best KRR model from Fig. 6 (cMBDF 82; 81), we performed a prediction error analysis to detect outliers within the dataset. We trained KRR/cMBDF based ML models on 200k molecules (4 disjointed training sets) to make predictions on the remaining dataset. This was done on all the 784,875 equilibrium geometries (including conformers) reported in the dataset. The resulting distribution of errors on the entire dataset is shown in fig. 7. For each molecule we used the prediction with the smallest error out of the 3 ML models that were not trained on it. The mean absolute error for the 784,875 predictions was obtained to be 0.75 kcal/mol with a standard deviation of 1.55 kcal/mol. The largest obtained error was 167.3 kcal/mol and the 25 largest outliers show a mean absolute error of 85.9 kcal/mol. These molecules are shown as insets in fig. 7.
Figure 7: Atomization energy prediction errors on the 784,875 converged molecules from the VQM24 dataset. All predictions were made by KRR models employing the cMBDF 81; 82 representation after training on 200k molecules. Errors for all molecules were obtained by training four such models with disjointed training sets and using the smallest prediction error for all out-of-sample molecules. Mean absolute error (MAE) for the predictions is shown in the top right inset along with the standard deviation. Insets show molecules with the largest prediction errors. Atomic colours correspond to: grey-Carbon, white-Hydrogen, blue-Nitrogen, red-Oxygen, dark green-Fluorine, cream-Silicon, orange-Phosphorous, yellow-Sulfur, green-Chlorine, dark red-Bromine.
Code Availability
Sample code for accessing the dataset is mentioned above. Code used for generating the data and more tools can be found at
https://github.com/dkhan42/VQM24